what are the chemical equations for all the experiment in anion analysis

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
below are the equations of anion analysis that u are supposed to form in the chemistry labrotory 
NACL+AGNO3 =AGCL+NANO3
CL-+AG+ = AGCL 
HCO3+H+ = CO2+H20 
CO2+BA+2OH- = BACO3 +H20 
BA(2+)+SO4(2-) = BASO4 
CL2+2I- = I2+2CL- 

i have not written the process as in the question u mentioned only reactions , if u need process , then just let me know or ask a question regarding that .. 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..